Mary Ott is going to borrow $14,600 for 45 days and pay $185 interest. What is the effective rate of interest if the loan is discounted? (Use a 360-day year. Do not round intermediate calculations. input your answer as a percent rounded to 2 decimal places.) Effective rate on a discounted loan I%
Answer #1

Effective rate on a discounted loan = Interest/(Prinicipal amount - Interest) * 360/loan days *100

= $185/($14,600-$185)*360/45*100

= $185/$14,415*8*100

= 10.267%
Answer #2

SOLUTION :


Loan is discounted.

Receipt on taking loan = (14600 - 185) = 14415 ($)

Interest of $185 is actually on $14415.


Effective annual rate

= [ ( 1 + interest rate of the period)^ (number of periods in a year) ] - 1

= (1 + 185/14415)^(360/45) - 1

= 0.1074

= 10.74 % (ANSWER).


answered by: Tulsiram Garg